# CADDesigner: Conceptual Design of CAD Models Based on General-Purpose Agent

### v2

@@ Abstract
Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing but typically requires a high level of expertise from designers. To lower the entry barrier and improve design efficiency, we present an agent for CAD conceptual design powered by large language models (LLMs). The agent accepts both abstract textual descriptions and freehand sketches as input, engaging in interactive dialogue with users to refine and clarify design requirements through comprehensive requirement analysis. Built upon a novel Context-Independent Imperative Paradigm (CIP), the agent generates high-quality CAD modeling code. During the generation process, the agent incorporates iterative visual feedback to improve model quality. Generated design cases are stored in a structured knowledge base, enabling continuous improvement of the agent’s code generation capabilities. Experimental results demonstrate that our method achieves state-of-the-art performance in CAD code generation.
@@ Introduction
Computer-Aided Design (CAD) software is widely used across production and manufacturing processes in a range of industries, including automotive, aerospace, and medical device sectors. Its adoption has significantly improved industrial productivity and design efficiency. Traditional CAD platforms—such as OnShape, AutoCAD, SolidWorks, and CATIA—typically rely on manual modeling by skilled designers with deep domain knowledge and technical expertise. However, recent advances in artificial intelligence, particularly the emergence of large language models (LLMs), present promising opportunities to automate CAD model generation, thereby reducing the entry barrier and accelerating the design workflow.
@@
Early research on automatic CAD model generation has primarily focused on parametric modeling approaches (Wu, Xiao, and Zheng 2021; Li et al. 2025c; Khan et al. 2024a, b). However, these methods are constrained by the limited diversity of training data and the representational capacity of their model outputs, supporting only a narrow set of CAD operations. Recent work leveraging large language models (LLMs) for CAD code generation has shown promise in overcoming these limitations (Li et al. 2025a; Wang et al. 2025; Rukhovich et al. 2024). Fine-tuned LLMs can interpret multimodal user inputs and generate corresponding CAD modeling code. Nevertheless, fine-tuning open-source LLMs requires substantial GPU resources, and high-quality CAD training datasets remain scarce, limiting the diversity and quality of the generated code.
@@
To address these challenges, we introduce CADDesigner, an LLM-powered agent for conceptual CAD design. The agent accepts both abstract textual descriptions and hand-drawn sketches as user input, engaging in interactive dialogues with designers to iteratively refine and clarify design requirements through comprehensive requirement analysis.
@@
To enhance the code generation process, we propose a novel Context-Independent Imperative Paradigm (CIP) that enables the agent to produce high-quality CAD modeling code. The agent leverages visual feedback throughout the generation process to incrementally improve model fidelity. Representative results generated by CADDesigner are shown in Figure 1, and our main contributions are summarized as follows:
@@
  • We present CADDesigner, a novel framework for CAD modeling code generation. CADDesigner accepts abstract textual descriptions and sketches as user input, and integrates tools for requirement analysis, knowledge-constrained code generation, and vision-based error correction to generate CAD scripts that satisfy user intent.
@@
  • We introduce a context-independent paradigm for CAD modeling script generation based on an imperative function-calling structure. This paradigm decouples individual CAD operations and improves code generation quality through explicit type annotations, LLM-friendly error representations, and self-evolving capabilities. It supports a wide range of operations, including extrusion, revolution, fillet, chamfer, sweeping, lofting, etc.
@@
Parametric CAD Model Generation
Parametric CAD model generation approaches commonly employ supervised learning techniques to produce sequences of CAD commands, which can be imported into CAD modeling software to generate editable, parametric files (Willis et al. 2021; Wu, Xiao, and Zheng 2021). Systems such as DeepCAD (Wu, Xiao, and Zheng 2021), Fusion 360 Gallery (Willis et al. 2021), SkexGen (Xu et al. 2022), and Diffusion-CAD (Zhang et al. 2025) adopt customized neural network architectures trained to generate command sequences. However, they primarily support only basic modeling operations such as sketching and extrusion.
Additional efforts (Khan et al. 2024a; Guo et al. 2022; Ma et al. 2024) explore CAD model reconstruction from point cloud data, while more recent methods (Li et al. 2025d; Chen et al. 2025) aim to infer structured, editable CAD models directly from unstructured RGB images (either single- or multi-view), bridging the gap between 2D visual understanding and 3D CAD modeling. Despite these advancements, existing approaches remain limited in their ability to generate complex CAD models, primarily due to constrained training data diversity and limited output representation capacity.
LLM-based CAD Model Generation
Fine-tuned large language models (LLMs) and vision-language models (VLMs) have emerged as a prominent direction for CAD model generation, enabling the synthesis of CAD command sequences or Python code from multimodal input. Cad-LLM (Wu et al. 2023b) and CadVLM (Wu et al. 2024) utilize fine-tuned LLMs to generate parametric models from sketches and images. CAD-MLLM (Xu et al. 2024) targets multimodal inputs—including text, images, and point clouds—for parametric model generation. CAD-Llama (Li et al. 2025a) proposes a hierarchical annotation pipeline that transforms command sequences into semantically structured CAD code. This is followed by adaptive pre-training and instruction tuning, enabling LLMs to generate high-fidelity parametric 3D models. CAD-Recode (Rukhovich et al. 2024) converts point clouds into sequential representations, then uses a pre-trained LLM (Qwen2-1.5B (Team 2024)) to generate corresponding CAD code. Text-to-CadQuery (Xie and Ju 2025) directly maps text descriptions to CadQuery code, leveraging pre-trained LLMs’ capabilities in Python generation and spatial reasoning.
@@
Agent-based CAD Model Generation
Another research direction focuses on constructing CAD generation agents using prompt engineering and ultra-large foundation models, circumventing the need for additional training. CAD-Assistant (Mallis et al. 2024) presents a general-purpose CAD agent framework that integrates visual and language models (VLLMs) as planners to generate Python code for FreeCAD, achieving zero-shot capability across diverse CAD tasks. SeekCAD (Li et al. 2025b) combines retrieval-augmented generation (RAG), visual feedback, and a chain-of-thought (CoT) mechanism to generate customized CAD modeling code with self-optimization. 3D-PreMise (Yuan et al. 2024) investigates the potential and limitations of LLMs in program synthesis for manipulating 3D software and generating parametrically controlled shapes. CADCodeVerify (Wu et al. 2023a) incorporates iterative validation and refinement loops, using visual language models to pose and answer verification questions about the generated CAD outputs.
@@
CAD Modeling Architecture
Figure 2: The CADDesigner agent follows a ReAct-style paradigm to progressively transform user requirements into valid CAD models through iterative reasoning, tool execution, and feedback refinement. It first refines user requirements into structured specifications, generates executable modeling code using domain APIs, and analyzes execution results via both symbolic (e.g., shell logs, geometric metadata) and visual feedback (e.g., rendered 3D views). In the figure, arrows of different colors indicate information flows across various stages of the modeling loop, while the numbered circles represent the temporal order of data generation. Readers are encouraged to follow the numbers sequentially to understand the agent’s internal workflow.
In the following subsections, we first present the architecture of the CADDesigner agent. We then provide an overview of the CAD tools used, followed by detailed descriptions of each individual tool integrated into the system.
ReAct Agent rather than Workflow
CADDesigner employs a ReAct-style (Yao et al. 2023) agent-centric architecture, where a central agent governs the entire modeling loop through iterative reasoning, tool execution, and feedback integration.
Unlike traditional hard-coded workflows, our architecture leverages prompt-based workflow specification, where system prompts encode procedural guidance for the agent. This design allows for flexible and revisable orchestration, as the workflow is not statically programmed but interpreted at runtime by the agent. As a result, users can intervene at any stage of the process, like adding constraints, correcting directions, or adjusting goals, leading to a truly interruptible and human-in-the-loop modeling cycle
@@
CAD Tools
@@
CADDesigner employs a set $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4}\}$ of four CAD-specific tools, each serving a distinct functional role in the design-to-modeling pipeline:
@@
  • $T_{1}:\mathcal{I}\to\mathcal{D}$ maps user-provided multimodal inputs $\mathcal{I}$ to detailed design descriptions $\mathcal{D}$, refining initial requirements into structured design specifications.
@@
  • $T_{2}:\mathcal{D}\to\mathcal{C}$ translates finalized design descriptions into executable CAD modeling code $\mathcal{C}$. A new code paradigm for CAD code generation has been designed and used by $T_{2}$. This will be discussed in detail in the following section “CAD Code Generation Paradigm”.
@@
  • $T_{3}:\mathcal{C}\to\mathcal{M}$ executes the generated code $\mathcal{C}$ to produce corresponding CAD models $\mathcal{M}$, aligning with standard shell command execution processes.
@@
  • $T_{4}:\{\mathcal{M},\mathcal{D}\}\to(P,F)$ returns a binary tuple where $P\in\{0,1\}$ indicates the termination flag ($P=1$ triggers termination), and $F$ is diagnostic feedback for potential refinement. Inside this tool, an multi-view render result of model $\mathcal{M}$ will be generated. We denote this render result as $\mathcal{V}$. Thus $T_{4}$ can be separated into 2 steps: $T_{4}^{1}:\mathcal{M}\to\mathcal{V}$, and $T_{4}^{2}:\{\mathcal{V},\mathcal{D}\}\to(P,F)$
@@
The sequential composition of these tools forms a functional pipeline:
$T_{4}^{1}\circ T_{3}\circ T_{2}\circ T_{1}:\mathcal{I}\to\mathcal{V}$
@@
This pipeline is augmented with a closed-loop correction mechanism: after generating visual feedback $\mathcal{V}$, the system invokes $T_{4}^{2}$ to obtain $(P,F)$. If $P=0$, the feedback $F$ is used to guide revised design generation via $T_{3}\circ T_{2}\circ T_{1}$. If $P=1$, the process terminates successfully. Thus, the loop continues until $T_{4}^{2}$ returns a termination signal.
Requirements Analysis
@@
During the conceptual design phase, designers typically have only rough descriptions or sketches, making it challenging to precisely define detailed model parameters as well as modeling process. To address this, we support designers by conducting requirement analysis to clarify and refine their design intention.
@@
Let the set of initial information input by the user be $\mathcal{I}(Text,Img)$, where TextText represents simple text description and ImgImg represents some sketchs. Let $\mathcal{P}$ be the set of necessary parameters required for generating a CAD model. The process of the CADDesigner conducting detailed design can be expressed as $F:\mathcal{I}\to\mathcal{D}_{\text{detail}}$, where $\mathcal{D}_{\text{detail}}$ is detailed designs containing specific parameters, satisfying $\mathcal{P}\subseteq\mathcal{D}_{\text{detail}}$, which is accomplished by $T_{1}$. The user’s revision process for the detailed design is an iterative function:
$R:\mathcal{D}_{\text{detail}}\times\mathcal{U}\to\mathcal{D}_{\text{detail}}$
@@
Knowledge Constrained Code Generation
@@
$\mathcal{D}_{\text{final}}$ is the detailed design confirmed by the user. The process by which the code generation tool $T_{2}$ receives $\mathcal{D}_{\text{final}}$ and generates CAD modeling code is expressed as $G:\mathcal{D}_{\text{final}}\to\mathcal{C}$. A knowledge base $K$ is constructed, which includes a set of function annotations $Anno\subseteq\mathcal{K}$ and a set of basic model cases $Case\subseteq\mathcal{K}$. The process by which the CADDesigner generates code with reference to the knowledge base can be expressed as
$G^{\prime}:\mathcal{D}_{\text{final}}\times\mathcal{K}\to\mathcal{C^{\prime}}$
and $\mathcal{C^{\prime}}\supset\mathcal{C}$, meaning that referring to the knowledge base can improve the success rate of code generation. The running result of the code is delivered to the judgment function $J:\mathcal{C}\to\{0,1\}$. When $J(\mathcal{C})=1$, the code can run successfully and generate a model $\mathcal{M}=T_{3}(\mathcal{C})$, which is then passed to the next stage; when $J(\mathcal{C})=0$, new code is regenerated based on the error feedback.
Visual Based Iterative Error Correction

### v4

@@ Abstract
Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing but typically requires a high level of expertise from designers. To lower the entry barrier and improve design efficiency, we present an agent for CAD conceptual design powered by large language models (LLMs). The agent accepts both textual descriptions and sketches as input, engaging in interactive dialogue with users to refine and clarify design requirements through comprehensive requirement analysis. Built upon a novel Explicit Context Imperative Paradigm (ECIP), the agent generates high-quality CAD modeling code. During the generation process, the agent incorporates iterative visual feedback to improve model quality. Generated design cases are stored in a structured knowledge base, enabling continuous improvement of the agent’s code generation capabilities. Experimental results demonstrate that our method achieves state-of-the-art performance in CAD code generation.
keywords:
[…]
@@ 1 Introduction
Computer-Aided Design (CAD) software is widely used across production and manufacturing processes in a range of industries, including automotive, aerospace, and medical device sectors. Its adoption has significantly improved industrial productivity and design efficiency. Traditional CAD platforms—such as OnShape, AutoCAD, SolidWorks, and CATIA—typically rely on manual modeling by skilled designers with deep domain knowledge and technical expertise. However, recent advances in artificial intelligence, particularly the emergence of LLMs, present promising opportunities to automate CAD model generation, thereby reducing the entry barrier and accelerating the design workflow.
@@
Early research on automatic CAD model generation has primarily focused on parametric modeling approaches [wu2021deepcad, li2025mamba, khan2024cad, text2cad]. However, these methods are constrained by the limited diversity of training data and the representational capacity of their model outputs, supporting only a narrow set of CAD operations. Recent work leveraging LLMs for CAD code generation has shown promise in overcoming these limitations [li2025cad, wang2025cad, rukhovich2024cad]. Fine-tuned LLMs can interpret multimodal user inputs and generate corresponding CAD modeling code. Nevertheless, fine-tuning open-source LLMs requires substantial GPU resources, and high-quality CAD training datasets remain scarce, limiting the diversity and quality of the generated code.
@@
To address these challenges, we introduce CADDesigner, an LLM-powered agent for conceptual CAD design. The agent accepts both textual descriptions and sketches as user input, engaging in interactive dialogues with designers to iteratively refine and clarify design requirements through comprehensive requirement analysis.
@@
To enhance the code generation process, we propose a novel Explicit Context Imperative Paradigm (ECIP) that enables the agent to produce high-quality CAD modeling code. The agent leverages visual feedback throughout the generation process to incrementally improve model fidelity. Representative results generated by CADDesigner are shown in Figure 1, and our main contributions are summarized as follows:
@@
  • We present CADDesigner, a novel framework for CAD modeling code generation. CADDesigner accepts textual descriptions and sketches as user input, and integrates tools for requirement analysis, knowledge-constrained code generation, and vision-based error correction to generate CAD scripts that satisfy user intent.
@@
  • We introduce an explicit context imperative paradigm for CAD modeling script generation based on an imperative function-calling structure. This paradigm decouples individual CAD operations and improves code generation quality through explicit type annotations, LLM-friendly error representations, and self-evolving capabilities. It supports a wide range of operations, including extrusion, revolution, fillet, chamfer, sweeping, lofting, etc.
@@ 2 Related Work
2.1 Parametric CAD Model Generation
Parametric CAD model generation approaches commonly employ supervised learning techniques to produce sequences of CAD commands, which can be imported into CAD modeling software to generate editable, parametric files [willis2020fusion, wu2021deepcad]. Systems such as DeepCAD [wu2021deepcad], Fusion 360 Gallery [willis2020fusion], SkexGen [xu2022skexgen], and Diffusion-CAD [zhang2025diffusion] adopt customized neural network architectures trained to generate command sequences. However, they primarily support only basic modeling operations such as sketching and extrusion.
Additional efforts [khan2024cad, guo2022complexgen, ma2024draw] explore CAD model reconstruction from point cloud data, while more recent methods [li2025caddreamer, chen2025cadcrafter] aim to infer structured, editable CAD models directly from unstructured RGB images (either single- or multi-view), bridging the gap between 2D visual understanding and 3D CAD modeling. Despite these advancements, existing approaches remain limited in their ability to generate complex CAD models, primarily due to constrained training data diversity and limited output representation capacity.
2.2 LLM-based CAD Model Generation
Fine-tuned LLMs and vision-language models (VLMs) have emerged as a prominent direction for CAD model generation, enabling the synthesis of CAD command sequences or Python code from multimodal input. Cad-LLM [wu2023cad] and CadVLM [wu2024cadvlm] utilize fine-tuned LLMs to generate parametric models from sketches and images. CAD-MLLM [xu2024cad] targets multimodal inputs—including text, images, and point clouds—for parametric model generation. CAD-Llama [li2025cad] proposes a hierarchical annotation pipeline that transforms command sequences into semantically structured CAD code. This is followed by adaptive pre-training and instruction tuning, enabling LLMs to generate high-fidelity parametric 3D models. CAD-Recode [rukhovich2024cad] converts point clouds into sequential representations, then uses a pre-trained LLM (Qwen2-1.5B [team2024qwen2]) to generate corresponding CAD code. Text-to-CadQuery [xie2025text] directly maps text descriptions to CadQuery code, leveraging pre-trained LLMs’ capabilities in Python generation and spatial reasoning.
@@ 2.2 LLM-based CAD Model Generation
2.3 Agent-based CAD Model Generation
Another research direction focuses on constructing CAD generation agents using prompt engineering and ultra-large foundation models, circumventing the need for additional training. CAD-Assistant [mallis2024cad] presents a general-purpose CAD agent framework that integrates visual and language models (VLLMs) as planners to generate Python code for FreeCAD, achieving zero-shot capability across diverse CAD tasks. SeekCAD [li2025seek] combines retrieval-augmented generation (RAG), visual feedback, and a chain-of-thought (CoT) mechanism to generate customized CAD modeling code with self-optimization. 3D-PreMise [yuan20243d] investigates the potential and limitations of LLMs in program synthesis for manipulating 3D software and generating parametrically controlled shapes. CADCodeVerify [wu2023visual] incorporates iterative validation and refinement loops, using visual language models to pose and answer verification questions about the generated CAD outputs.
@@ 2.3 Agent-based CAD Model Generation
3 CAD Modeling Architecture
Figure 2: The Intelligent CAD Orchestrator Agent, CADDesigner, follows a ReAct-style paradigm to progressively transform user requirements into valid CAD models through iterative reasoning, tool execution, and feedback refinement. It first refines user requirements into detailed designs, generates executable modeling code using domain APIs, and analyzes execution results via both symbolic (e.g., shell logs and errors) and visual feedback (e.g., rendered 3D views).
In this section, we introduce the CADDesigner architecture, which features a ReAct-style agent interacting with a diverse set of integrated CAD tools to iteratively generate and refine CAD models. The overall framework is illustrated in Figure 2.
3.1 ReAct Agent rather than Workflow
CADDesigner employs a ReAct-style [yao2023react] agent-centric architecture, where a central agent governs the entire modeling loop through iterative reasoning, tool execution, and feedback integration.
Unlike traditional hard-coded workflows, our architecture leverages prompt-based workflow specifications, where the system prompts encode procedural guidance for the agent. This design allows for flexible and revisable task orchestration, as the workflow is not statically programmed but interpreted at runtime by the agent. As a result, users can intervene at any stage of the process, like adding constraints, correcting directions, or adjusting goals, leading to a truly interruptible and human-in-the-loop modeling cycle.
@@ 3.1 ReAct Agent rather than Workflow
3.2 CAD Tools
The intelligent orchestration relies on a diverse suite of CAD-oriented tools. Each tool addresses a specific stage within the design-to-modeling pipeline and works in synergy under the direction of the ReAct agent. This subsection describes the CAD tools integrated in our pipeline.
3.2.1 Core CAD Toolset
@@
CADDesigner employs various CAD-specific tools to support its design-to-modeling pipeline. Central to this set are four tools, denoted $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4}\}$, each fulfilling a distinct role.
@@
  • Requirement Refinement Tool $T_{1}\colon\mathcal{I}\to\mathcal{D}$ maps user-provided multimodal inputs $\mathcal{I}$ to detailed design descriptions $\mathcal{D}$, refining initial requirements into structured design specifications.
@@
  • Code Generation Tool $T_{2}\colon\mathcal{D}\to\mathcal{C}$ translates finalized design descriptions into executable CAD modeling code $\mathcal{C}$. A new code paradigm for CAD code generation has been designed and used by $T_{2}$. This will be discussed in detail in the section 4.
@@
  • Command Execution Tool $T_{3}\colon\mathcal{C}\to\mathcal{M}$ executes the generated Python code $\mathcal{C}$ to produce the corresponding CAD model $\mathcal{M}$ by running shell commands.
@@
  • Visual Feedback Tool $T_{4}\colon\{\mathcal{M},\mathcal{D}\}\to(P,F)$ returns a binary pair where $P\in\{0,1\}$ indicates the termination flag ($P=1$ triggers termination), and $F$ is diagnostic feedback for potential refinement. Inside this tool, a multi-view render result of model $\mathcal{M}$ will be generated. We denote this render result as $\mathcal{V}$. Thus $T_{4}$ can be separated into 2 steps: $T_{4}^{1}:\mathcal{M}\to\mathcal{V}$, and $T_{4}^{2}:\{\mathcal{V},\mathcal{D}\}\to(P,F)$
@@
The sequential composition of these core tools forms the main functional pipeline:
$T_{4}^{1}\circ T_{3}\circ T_{2}\circ T_{1}:\mathcal{I}\to\mathcal{V}.$ (1)
@@
This pipeline is augmented with a closed-loop correction mechanism: after generating visual feedback $\mathcal{V}$, the system invokes $T_{4}^{2}$ to obtain $(P,F)$. If $P=0$, the feedback $F$ is used to guide revised design generation via $T_{3}\circ T_{2}\circ T_{1}$. If $P=1$, the process terminates successfully. Otherwise, the loop continues until $T_{4}^{2}$ returns a termination signal.
3.2.2 Auxiliary Toolset
In addition to the core CAD toolset, CADDesigner integrates several auxiliary tools that support efficient context management and system interaction. These include:
  • SketchPad Tool: A flexible key-value storage tool designed to mitigate the context explosion problem in LLM-based agents. SketchPad allows the agent to store arbitrary intermediate data at any time during the design process, including image paths, reference code snippets, execution results, and more. By storing context data as keyed entries instead of embedding entire data repeatedly, it significantly reduces LLM context size. SketchPad also supports automatic summarization, tag-based search, and automatic expiration of outdated entries, enabling flexible and scalable context management.
  • File Operation Tools: These tools provide fundamental file-reading and file-writing capabilities, allowing the agent to store or load files as required throughout the CAD design pipeline. Similar to SketchPad, file content or references are stored and accessed efficiently to avoid large context overheads.
Together, these auxiliary tools complement the core CAD toolset by enabling robust, efficient, and scalable operation of the CADDesigner agent in complex multi-turn interactions.
3.3 Requirements Analysis
@@
During the conceptual design phase, designers typically have only rough descriptions or sketches, making it challenging to precisely define detailed model parameters as well as the modeling process. To address this, we support designers by conducting requirement analysis to clarify and refine their design intention.
@@
Let the set of initial information input by the user be $\mathcal{I}\subseteq\mathcal{I}_{\mathrm{txt}}\times\mathcal{I}_{\mathrm{img}}$, where $\mathcal{I}_{\mathrm{txt}}$ represents the set of possible text inputs (including empty text $\emptyset$) and $\mathcal{I}_{\mathrm{img}}$ represents the set of possible image inputs (including a null image $\emptyset$). Let $\mathcal{P}$ be the set of necessary parameters required for generating a CAD model. The process of the CADDesigner conducting detailed design can be expressed as $\mathcal{F}_{\text{design}}:\mathcal{I}\to\mathcal{D}_{\text{detail}}$, where $\mathcal{D}_{\text{detail}}$ is detailed designs containing specific parameters, satisfying $\mathcal{P}\subseteq\mathcal{D}_{\text{detail}}$, which is accomplished by $T_{1}$. The user’s revision process for the detailed design is an iterative function:
$R:\mathcal{D}_{\text{detail}}\times\mathcal{U}\to\mathcal{D}_{\text{detail}},$ (2)
@@
3.4 Knowledge Constrained Code Generation
@@
$\mathcal{D}_{\text{final}}$ is the detailed design confirmed by the user. The process by which the code generation tool $T_{2}$ receives $\mathcal{D}_{\text{final}}$ and generates CAD modeling code is expressed as $G:\mathcal{D}_{\text{final}}\to\mathcal{C}$. A knowledge base $\mathcal{K}$ is constructed to support code generation. $\mathcal{K}$ includes a set of function annotations $Anno\subseteq\mathcal{K}$ and a set of basic model cases $Case\subseteq\mathcal{K}$, which provide semantic and structural references for the code generation tool. With the assistance of the knowledge base, the code generation process is extended to:
$G^{\prime}:\mathcal{D}_{\text{final}}\times\mathcal{K}\to\mathcal{C}^{\prime},$ (3)
where $\mathcal{C}^{\prime}$ denotes the code generated referencing the knowledge base, which is semantically and structurally richer compared to the code $\mathcal{C}$ generated without using the knowledge base. The execution result of the code is delivered to the judgment function $J:\mathcal{C}^{\prime}\to\{0,1\}$. When $J(\mathcal{C}^{\prime})=1$, the code can run successfully and generate a model $\mathcal{M}=T_{3}(\mathcal{C}^{\prime})$, which is then passed to the next stage; when $J(\mathcal{C}^{\prime})=0$, the error information is fed back along with the knowledge base $\mathcal{K}$ to guide the next iteration of code generation $G^{\prime}$ to produce a new code. This feedback loop helps to iteratively correct the code and improve the success rate. Therefore, the knowledge base provides essential semantic and structural support for code generation, and the iterative process combined with execution feedback further enhances the likelihood of successful code generation.
3.5 Visual-based Iterative Error Correction